# Visualising Policy-Reward Interplay to Inform Zeroth-Order Preference Optimisation of Large Language Models

### v1

@@ 6 Conclusions
In this work, we presented a comprehensive visual and quantitative analysis of Preference Optimisation dynamics by tracking the evolution of both policy and reward models through popular methods. Our visualisations using t-SNE and UMAP revealed structured training trajectories, within each model and in relation to each other. These insights guided the development of ZOPrO, a novel Zeroth-Order Optimisation method that uses our empirical findings to restrict its gradient search and thus improve convergence rate.
ZOPrO implements a modified SPSA approach with an adaptive perturbation strategy that balances a deterministic component (from prior policy updates) and restricted stochastic exploration. We show how this choice preserves theoretical convergence guarantees. Our experiments on the tasks of summarisation, machine translation and conversational assistants, show that ZOPrO achieves consistent improvements in reward while maintaining a lower memory footprint and faster per-iteration runtime than first-order methods. This inherently comes with less improvement per step, putting our method just behind state-of-the-art performance. Nevertheless, we demonstrate that gradient-free optimisation can be both viable and competitive even when targeting Preference Optimisation.
7 Limitations
Despite the promising results, several limitations warrant discussion. First, while ZOPrO exhibits faster training iterations and lower memory usage, its per-step improvements are inherently noisier compared to first-order methods like RLOO. Further, it necessitates careful tuning of hyper-parameters such as the perturbation multiplier $\varepsilon$ and the learning rate $\eta$. During our testing, we uncovered that similar to other ZO methods, ours has an over-sensitivity to these settings as well. Future work could explore more advanced scheduling or tuning of these parameters for more robust results.
@@ 7 Limitations
Finally, our experiments were conducted on a relatively small model (LLaMA-3.2 1B) but varying tasks and domains (conversation, summarisation, and machine translation). As such, the scalability of ZOPrO to larger models remains to be fully evaluated.
@@ Appendix B Training details
All experiments, unless otherwise specified, were conduced on 4$\times$×A100 with 40GB. As the model can fit into a single GPU, the parallelism was used to speed up training. All experiments were also conducted in half precision, bfloat16. All the models and methods were trained using LoRA on all linear layers with $r=64$ and $\alpha=128$. We set the maximum sequence length to 1024 and the maximum generation length at 53. For all the methods except ours, we use a learning rate of $55e-$ 5 which linearly decays. For our method, we still use a linear decay but explore values in $[1e-3,5e-5,1e-4,5e-5,1e-5]$. For our method, we also explore $\varepsilon\in[1e-2,1e-3,1e-4,1e-5,1e-6]$. The best combination was $\varepsilon=1e-5$ and $\eta=1e-4$. For all methods experiments we keep a fixed batch size of 64. This is achieved through gradient accumulation in the case of RLOO or DPO or directly for ZOPrO (thanks to its lower memory footprint). A single iteration takes 1h20m for ZOPrO, 1h49m for DPO, 5h for RLOO, 5h30 for PPO.
@@ B.2 Data details
HH-RLHF dataset contains 161K samples in the training split and 8k in the test split. All samples are in English. Summarize from Feedback has 93k samples in the training split and 86k in the validation. WMT has 7k samples for training and 1k for validation. For WMT we only use the en-de split, as that is the only language supported by Llama 3.2.

### v2

@@ 6 Conclusions
In this work, we presented a comprehensive visual and quantitative analysis of Preference Optimisation dynamics by tracking the training evolution of both policy and reward models through popular methods. Our visualisations using t-SNE and UMAP revealed structured training trajectories, within each model (e.g., policy at iteration $t$ compared to $t+1$) and in relation to each other (i.e., policy compared to reward). These insights guided the development of ZOPrO, the first Zeroth-Order method applied to the task of Preference Optimisation, where all previous methods only considered classification or question-answering.
ZOPrO uses our empirical findings to restrict its gradient search and thus improves its convergence rate. This is done by implementing a modified SPSA approach with an adaptive perturbation strategy that balances a deterministic component (from prior policy and reward updates) and restricted stochastic exploration. We show how this choice preserves theoretical convergence guarantees. Our experiments on the tasks of summarisation, machine translation and conversational assistants, show that ZOPrO achieves consistent improvements in reward while maintaining a lower memory footprint and faster per-iteration runtime than first-order methods. This inherently comes with less overall improvement per step, putting our method just behind state-of-the-art performance. Nevertheless, we demonstrate that gradient-free, Zeroth-Order optimisation can be both viable and competitive even when targeting Preference Optimisation.
Acknowledgements
We thank Caterina Nottolini for her valuable help in double-checking the mathematical details and for insightful brainstorming throughout the development of this work.
The computations and data handling were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Alvis, C3SE (Chalmers) partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
Limitations
Despite the promising results, several limitations warrant discussion. First, while ZOPrO exhibits faster training iterations and lower memory usage, its per-step improvements are inherently noisier compared to first-order methods like RLOO. Further, it necessitates careful tuning of hyperparameters such as the perturbation multiplier $\varepsilon$ and the learning rate $\eta$. During our testing, we did a sweep over 9 combinations of these hyperparameters. Whilst almost all combinations avoided model collapse and more than half showed significant improvement, future work could explore more advanced scheduling or tuning of these hyperparameters for more robust results.
@@ Acknowledgements
ZOPrO was implemented following RLOO’s objective. Whilst generalising it to similar objectives (e.g., PPO) can be straightforward, our method can only be applied to iterative training regimes. Indeed, the relaxation of the sampling strategy could allow the adaptation to other objectives (such as standard DPO), however, this would cause a slowdown in the convergence rate. We leave such investigation to future works.
Our experiments were conducted on relatively small models (1.2-2.4B) but varying tasks (conversation, summarisation, and machine translation) and model families (Llama, Qwen and Gemma). As such, the scalability of ZOPrO to larger models remains to be fully evaluated.
@@ Appendix B Training details
All experiments, unless otherwise specified, were conducted on 4$\times$×A100 with 40GB of VRAM. As the model can fit into a single GPU, the parallelism was used to speed up training with DDP. All experiments were conducted in half precision, bfloat16. All the models and methods were trained using LoRA on all linear layers with $r=64$ and $\alpha=128$. We set the maximum sequence length to 1024 and the maximum generation length to 53. For all the methods except ours, we use a learning rate of $55e-$ 5 which linearly decays. For our method, we still use a linear decay but explore values in $[1e-3,1e-4,1e-5]$. For our method, we also explore $\varepsilon\in[1e-3,1e-4,1e-5]$. The best combination was $\varepsilon=1e-4$ and $\eta=1e-5$ for the conversation task. For all methods and experiments, we keep a fixed batch size of 64. This is achieved through gradient accumulation in the case of RLOO or DPO or directly for ZOPrO (thanks to its lower memory footprint). A single iteration (for the conversation task) takes 1h20m for ZOPrO, 1h49m for DPO, 5h for RLOO, 5h30 for PPO.
@@ B.2 Data details
HH-RLHF dataset contains 161K samples in the training split and 8k in the test split. All samples are in English. Summarize from Feedback has 93k samples in the training split and 86k in the validation. WMT has 7k samples for training and 1k for validation. For WMT we only use the en-de split.
Appendix C Example generations
We show in Table 11 some example generations from Llama 3.2 after training with our method for two iterations.